OpenStudy (vortish):

use the multiplication property of inequalities to sloe the inequality 5/4 >or equal to -3/8

OpenStudy (anonymous):

I'm not sure I know exactly what they're thinking for this one. I might need to look up that property first. It seems obvious by looking at it, so they clearly want you to do some multiplying to prove (or "show") that it is true.

OpenStudy (vortish):

that is right they want the solution set then look at the graphs at the bottom and find the right one

OpenStudy (vortish):

do you have skype jake

OpenStudy (anonymous):

could be something like: 5/4 >= -3/8 multiply both sides by 8: 8(5/4) >= 8(-3/8) simplify: 40/4 >= -24/8 reduce: 10 >= -3 (true... obviously) I don't know... don't really see the point in going through these steps to show that a positive number is greater than or equal to a negative number.

OpenStudy (anonymous):

no skype, sorry.

OpenStudy (anonymous):

what sort of graphs are they? Number lines?

OpenStudy (vortish):

k they have a feature that allows you to see what is on my screen

OpenStudy (anonymous):

you can take a screen shot with ALT + PRT SCR on a windows computer... then paste it into a jpg file or a Word doc or whatever... you can upload the picture.

OpenStudy (vortish):

yea i do that so you can see what the heck im work with

OpenStudy (anonymous):

it's: 5/4 x >= 3/8 you left the "x" out in this question... changes things :)

OpenStudy (anonymous):

Huh? OS reorganized my posts... anyway, 5/4 x >= - 3/8 So multiply both sides by 4/5 to get rid of everything but the "x" on the left...

OpenStudy (vortish):

sorry about that

OpenStudy (vortish):

what I come up with is 10x >= -32/15

OpenStudy (anonymous):

so: (4/5)(5/4)x >= (-3/8)(4/5) simplify: x >= -12/40 reduce fraction on right: x >= -3/10

OpenStudy (vortish):

three will not go in to ten

OpenStudy (anonymous):

That's ok, you don't need it to divide... it's just like 0.3 or 30% or whatever... just a number that x must be greater than or equal to.

OpenStudy (vortish):

so the solution is x>= 3/10 and then the line would go to the left right

OpenStudy (anonymous):

lemme check my work... that answers not on your list.

OpenStudy (anonymous):

oops, we dropped the negative... it's x >= - 3/10 or x >= -0.3

OpenStudy (anonymous):

It starts at x = -0.3, and runs to the right for all points greater than -0.3

OpenStudy (anonymous):

I didn't really follow how you might have gotten a different result above, but if you follow what I did with simplifying the inequality, that's all there is.

OpenStudy (anonymous):

Do you know the key thing with multiplying across inequalities... the deal with multiplying or dividing by a negative number?

OpenStudy (anonymous):

Like if you have -2x >= 4.... you need to divide both sides by -2 to solve for the x term...

OpenStudy (anonymous):

But when you divide by -2, you have to reverse the direction of the inequality... -2x >= 4 but (-2)/(-2) <= 4/(-2) See how it flipped directions?